# Random Forests as Statistical Procedures: Design, Variance, and Dependence

### v2

@@ 4.4 The covariance floor and its strict positivity
Sections 4.2 and 4.3 identified two distinct mechanisms contributing to $C_{T}(x)$: reuse of training observations in local averages at $x$, and alignment of the induced local averaging rules across trees. We now show that observation reuse alone is sufficient to guarantee strict positivity of the covariance floor, without invoking any alignment assumptions.
@@
Theorem 2 shows that repeated reuse of the same training outcomes in local averages at $x$ is sufficient to induce a strictly positive covariance floor. This effect is governed entirely by the tree-generating design acting on the fixed covariate configuration and persists regardless of the number of trees aggregated. The condition $\mathbb{P}\!\bigl(W_{i^{\star}}(x;\theta)>0\mid X\bigr)>0$ is satisfied by all standard implementations whenever $x$ lies in the support of the design and at least one observation in its neighborhood has positive conditional variance; moreover, the alignment component provides a second, independent nonnegative contribution to the covariance floor that is not required for strict positivity but reinforces it.
4.5 Outcome stability of tree-based predictors
The preceding sections establish that random forest predictors exhibit irreducible dependence across trees induced by the design of the tree-generating mechanism. We now establish a complementary structural property of tree-based predictors: conditional on a realized tree structure, predictions at a fixed point are uniformly stable with respect to perturbations of the observed outcomes. This result concerns sensitivity to outcome perturbations holding the induced partition fixed, and is therefore conceptually distinct from the design-induced dependence analyzed above.
Fix a realized covariate configuration $X=\{X_{i}\}_{i=1}^{n}$ and a prediction point $x$. Conditional on a realized tree structure (that is, conditional on $X$ and a fixed draw of the tree-generating mechanism), the induced predictor at $x$ can be written in weighted form as
$\tilde{f}(x;Y)=\sum_{i=1}^{n}W_{i}(x)\,Y_{i},$
where the weights satisfy $W_{i}(x)\geq 0$ and $\sum_{i=1}^{n}W_{i}(x)=1$.
In particular, conditional on the realized tree structure, the mapping $Y\mapsto\tilde{f}(x;Y)$ is $1$-Lipschitz with respect to the $\ell_{\infty}$ norm.
Proof.
By linearity,
$\tilde{f}(x;Y+\delta Y)-\tilde{f}(x;Y)=\sum_{i=1}^{n}W_{i}(x)\,\delta Y_{i}.$
Because the weights are nonnegative and sum to one,
Remark 2 (Contrast with coefficient-based predictors).
Proposition 1 highlights a structural distinction between tree-based predictors and coefficient-based procedures. For linear and generalized linear models, predictions at a point $x$ depend on the outcomes through fitted coefficients involving an inverse information matrix, so small perturbations in $Y$ can be amplified under near collinearity, with sensitivity governed by the conditioning of the design matrix. By contrast, conditional on their realized structure, tree-based predictors depend on the outcomes only through nonnegative weights summing to one, yielding a uniform $\ell_{\infty}$ stability bound that is independent of predictor collinearity or design matrix conditioning. Instability in tree-based methods therefore arises through randomness in the induced partition, not through amplification of outcome noise.
@@ 5.1 The infinite-aggregation target and resolution
Designs that assign non-negligible expected weight to many observations are said to have lower resolution at $x$, while designs that concentrate expected weight on fewer observations localized near $x$ are said to have higher resolution at $x$.
@@
We use the term resolution because it more precisely characterizes the fineness of the forest’s local averaging rule and does not correspond to bias in the conventional parametric sense. Resolution describes the effective granularity of the infinite-aggregation predictor at $x$ under fixed $X$. Coarser weight distributions average responses over larger sets of observations, yielding increased stability across tree realizations but reduced sensitivity to local variation in the observed outcomes. Finer weight distributions concentrate mass on fewer observations, producing predictors that adapt more sharply to local structure in the realized data but that are more susceptible to dependence across trees through alignment of the induced averaging rules. All forest design parameters affect resolution through their influence on the distribution of the random weight vector $W(x;\theta)$.
@@ 5.2 Training fraction and resolution scaling
A defining feature of random forest constructions is that each tree is trained on a strict subset of the available observations. As a consequence, even under infinite aggregation the induced predictor reflects the resolution attainable by trees grown on a reduced training size. Under bootstrap resampling, the expected number of distinct observations used to construct a tree is approximately $0.632n$ (Efron, 1983). Under subsampling without replacement, the per-tree training size is fixed at $m=p_{\mathrm{obs}}\,n$ for some training fraction $p_{\mathrm{obs}}<1$. In both cases, the distribution of the weight vector $W(x;\theta)$ depends on $m$, and therefore the resolution of $f_{\infty}(x)$ is determined by the per-tree training size rather than by the full sample size.
Although a random forest reuses observations across the ensemble, each individual tree is constructed using only $m<n$ observations. The infinite-aggregation predictor $f_{\infty}(x)$ is therefore an expectation over tree-level prediction rules induced by trees grown on subsamples of size $m$, not a prediction rule trained on all $n$ observations. Averaging across trees combines these subsample-based predictors but does not increase the resolution beyond that attainable by the underlying tree-level fits.
@@
Whenever the distribution of $W(x;\theta)$ varies with the per-tree training size, $f_{\infty,m}(x)$ differs from $f_{\infty,n}(x)$ for $m<n$, and this discrepancy persists as the number of trees tends to infinity. Thus, reducing the per-tree training size induces a structural reduction in resolution that cannot be compensated for by aggregation.
@@ 5.3 Structural error pathways
The strict positivity of the covariance floor established in Theorem 2 implies that, even under infinite aggregation, the forest predictor retains design-induced variability determined by the distribution of the weight vector $W(x;\theta)$. While Monte Carlo variability becomes negligible with increasing $B$, the structural component $C_{T}(x)$ reflects how the randomized design allocates weight across training observations and how these allocations co-vary across trees.
Resolution and dependence arise from different aspects of the same design-induced randomness under fixed $X$. The infinite-aggregation predictor $f_{\infty}(x)=\mathbb{E}[T_{\theta}(x)\mid X]$ is determined by the average weight vector $\mathbb{E}[W(x;\theta)\mid X]$, which governs the resolution of the forest at $x$. In contrast, the covariance floor $C_{T}(x)$ depends on the variability and co-movement of the random weights across design realizations. Any design choice that reduces dependence necessarily alters the distribution of these weight vectors, and therefore also affects the resolution of the infinite-aggregation predictor. This relationship is analogous to the classical bias–variance tradeoff. Structural error pathways are thus governed by how the forest design redistributes variability between resolution and dependence, rather than by Monte Carlo variance reduction alone.
@@ 6.1 Parameterizing the tree-generating mechanism
We parameterize the tree-generating mechanism $\theta$ through a collection of design components defined relative to the observed sample size $n$. Although $n$ is not a tuning parameter in the same sense as subsampling or split restriction, it plays a structural role by determining the support over which the induced averaging sets $A_{\theta}(x)$ and weight vectors $W(x;\theta)$ are defined, and therefore influences both resolution and dependence.
@@
    Observation selection. A random index set $\mathcal{S}_{\theta}\subset\{1,\dots,n\}$ is drawn with fixed cardinality
    $m=|\mathcal{S}_{\theta}|,\qquad p_{\mathrm{obs}}=m/n,$
    defining the per-tree observation-selection fraction. The corresponding indicator vector $\mathcal{I}_{\theta}=(I_{\theta,1},\dots,I_{\theta,n})$, where $I_{\theta,i}:=\mathbf{1}\{i\in\mathcal{S}_{\theta}\}$, encodes which observations are exposed to the tree construction and governs eligibility for inclusion in the induced averaging sets $A_{\theta}(x)$.
@@
    Split-candidate restriction. At each internal node, a random candidate set of variables $\mathcal{J}_{\theta,v}\subset\{1,\dots,p\}$ is drawn with fixed size $q$. The split at that node is chosen by optimizing the splitting criterion over $\mathcal{J}_{\theta,v}$. This mechanism governs variability in the induced partitioning map $M_{\theta}$ and therefore controls alignment of local averaging rules across trees.
@@
    Terminal occupancy constraint. A stopping rule enforces a lower bound $s\geq 1$ on the size of the induced averaging set,
    $|A_{\theta}(x)|\geq s\quad\text{almost surely for all }x\text{ under consideration},$
    ensuring that each tree-level prediction averages at least $s$ observed outcomes. This constraint directly controls the sparsity and concentration of the weight vectors $W(x;\theta)$ and hence the effective resolution of the induced predictor.
@@
    Aggregation level. The forest predictor averages $B$ independent realizations $\theta_{1},\dots,\theta_{B}$. This parameter governs Monte Carlo approximation error but does not affect resolution or dependence in the infinite-aggregation limit.
@@
The forest design is therefore characterized by the parameter vector $(n,p_{\mathrm{obs}},q,s,B)$. For fixed $n$, the parameters $(p_{\mathrm{obs}},q,s)$ jointly determine the distribution of the induced averaging sets $A_{\theta}(x)$ and weight vectors $W(x;\theta)$, and thereby control both the resolution of the infinite-aggregation target and the dependence structure across trees. The role of $B$ is purely to approximate this design-based target through Monte Carlo averaging.
6.2 Structural error pathways and design parameters
Sections 3 and 5 characterize how the forest design induces distinct sources of error in the finite forest predictor at a fixed prediction point $x$. Finite aggregation introduces Monte Carlo variability around the infinite-aggregation target, while structural dependence across trees induces a variance component that persists as the number of trees increases. The covariance decomposition in Section 4 further separates this dependence into contributions arising from shared training observations and from alignment of the induced local averaging rules.
The design parameters $(p_{\mathrm{obs}},q,s,B)$ act on these error pathways through distinct and interpretable mechanisms.
  1. 1.
    Aggregation level ($B$). The number of trees controls only the Monte Carlo component arising from finite aggregation. Increasing $B$ improves approximation of the infinite-aggregation target $f_{\infty}(x)$ but does not alter the induced averaging sets $A_{\theta}(x)$, the distribution of weights $W(x;\theta)$, or the dependence structure across trees. Consequently, $B$ has no effect on resolution or on the design-induced covariance floor.
  2. 2.
    Training fraction ($p_{\mathrm{obs}}$). The training fraction governs which observations are exposed to each tree through the resampling indicators $\mathcal{I}_{\theta}$. Increasing $p_{\mathrm{obs}}$ increases joint inclusion probabilities and therefore strengthens the observation-reuse component of the covariance. At the same time, it expands the support of the induced averaging sets $A_{\theta}(x)$, allowing finer localization and higher resolution in the infinite-aggregation target. Decreasing $p_{\mathrm{obs}}$ weakens dependence at the cost of coarser local averaging.
  3. 3.
    Candidate-set size ($q$). The number of candidate variables considered at each split governs variability in the induced partitioning map $M_{\theta}$. Larger values of $q$ increase the likelihood that similar splitting decisions are made across trees, strengthening alignment of the induced local averaging rules and increasing the alignment component of the covariance. Smaller values of $q$ introduce additional randomness into $M_{\theta}$, weakening alignment but simultaneously broadening the distribution of induced averaging sets and reducing resolution.
  4. 4.
    Terminal occupancy constraint ($s$). The minimum size constraint on the averaging sets $A_{\theta}(x)$ directly controls how concentrated the weight vectors $W(x;\theta)$ can be. Larger values of $s$ enforce coarser local averages with greater stability but reduced resolution, while smaller values permit tighter localization around $x$ at the cost of increased single-tree variability and stronger dependence across trees.
Together, these parameters determine how the forest design reallocates error between Monte Carlo variability, structural dependence, and resolution of the induced predictor. No choice of design parameters can simultaneously minimize all sources of error: reductions in dependence are achieved only by altering the distribution of induced averaging sets, and gains in resolution necessarily increase sensitivity to design-induced dependence.
6.3 Aggregation level $B$
The aggregation level $B$ controls the number of independently generated trees averaged to form the forest predictor. Unlike the other design parameters, $B$ does not influence how individual trees are constructed or how predictions are formed within each tree. Instead, it governs only the degree of Monte Carlo averaging applied to a fixed randomized tree-generating mechanism.
Because the forest predictor is an average of $B$ independent realizations of the same randomized tree-level prediction rule, varying $B$ does not alter the distribution of the underlying tree predictor $T_{\theta}(x)$. In particular, aggregation does not affect the induced averaging sets $A_{\theta}(x)$, the weight distributions $W(x;\theta)$, or any dependence arising from repeated reuse of the observed data across trees.
Remark 4 (Aggregation controls Monte Carlo variability).
For any fixed tree-generating design and realized covariate configuration $X$,
$\mathbb{E}\!\left[\hat{f}_{B}(x)\mid X\right]=f_{\infty}(x),\qquad\operatorname{Var}\!\left(\hat{f}_{B}(x)\mid X\right)=\frac{1}{B}\,\sigma_{T}^{2}(x)+\frac{B-1}{B}\,C_{T}(x).$
Varying $B$ therefore affects only the Monte Carlo variability arising from finite aggregation. The infinite-aggregation target $f_{\infty}(x)$, the single-tree dispersion $\sigma_{T}^{2}(x)$, and the design-induced covariance $C_{T}(x)$ are properties of the tree-generating mechanism itself under fixed $X$ and are invariant to $B$.
This distinction is essential for interpreting the behavior of random forests. Increasing $B$ suppresses Monte Carlo noise around the design-induced target but cannot modify the resolution of the predictor or reduce dependence induced by the tree-generating mechanism. In particular, aggregation alone cannot eliminate the variance floor identified in Section 4.4.
6.4 Observation-selection fraction $p_{\mathrm{obs}}$
The observation-selection fraction $p_{\mathrm{obs}}$ is a fundamental design parameter governing how much of the observed dataset is exposed to each tree. Unlike the aggregation level $B$, which affects only Monte Carlo averaging, $p_{\mathrm{obs}}$ directly alters both the dependence structure across trees and the resolution of the induced predictor.
@@ 6.4 Observation-selection fraction $p_{\mathrm{obs}}$
From the perspective of dependence, $p_{\mathrm{obs}}$ acts exclusively through the observation-overlap component of the covariance decomposition introduced in Section 4.1 and analyzed in Section 4.2. In that representation, $p_{\mathrm{obs}}$ enters only through joint inclusion probabilities
@@
which determine how frequently the same observation can receive positive weight at $x$ in multiple trees. Increasing $p_{\mathrm{obs}}$ increases these joint inclusion probabilities and, holding the within-tree partitioning mechanism fixed, strengthens the observation-overlap contribution to $C_{T}(x)$.
Remark 5 (Monotonicity of observation-overlap dependence).
Holding the partitioning mechanism fixed, the observation-overlap contribution $C_{\mathrm{obs}}(x)$ is monotone increasing in the joint inclusion probabilities $\mathbb{P}(I_{\theta,i}=1,\;I_{\theta^{\prime},i}=1)$. This follows directly from the representation in Section 4.2, where these probabilities appear as nonnegative multiplicative factors.
Independently of its effect on dependence, $p_{\mathrm{obs}}$ governs resolution by restricting which observations are available to define candidate splits and terminal regions. Changing the observation-selection fraction alters the induced distribution of averaging sets $A_{\theta}(x)$, equivalently the induced distribution of weights $W(x;\theta)$, and therefore changes the infinite-aggregation target $f_{\infty}(x)$. Smaller values of $p_{\mathrm{obs}}$ induce coarser effective averaging by limiting the amount of data available to each tree, while larger values increase resolution by averaging over prediction rules trained on larger subsets of the observed data, at the cost of stronger reuse-driven dependence across trees.
Subsampling without replacement makes this trade-off explicit. When each tree is grown on a subset of size $m=p_{\mathrm{obs}}\,n$ drawn without replacement, $p_{\mathrm{obs}}$ becomes a tunable design parameter that directly controls both resolution and dependence. Decreasing $p_{\mathrm{obs}}$ weakens observation reuse across trees but reduces the resolution attainable by the induced predictor; increasing $p_{\mathrm{obs}}$ improves resolution while strengthening dependence through repeated reuse of the same observations.
The bootstrap corresponds to a particular, non-tunable choice of $p_{\mathrm{obs}}$. Although each tree is grown from a resample of size $n$, the expected fraction of distinct observations exposed to the tree satisfies
$\mathbb{P}(i\in\mathcal{I}_{\theta})=1-(1-1/n)^{n}\approx 1-e^{-1},\qquad\mathbb{P}(i\in\mathcal{I}_{\theta}\cap\mathcal{I}_{\theta^{\prime}})\approx(1-e^{-1})^{2}.$
Thus, under bootstrap resampling the observation-selection fraction is implicitly fixed by the resampling scheme, and the resulting magnitude of observation-induced dependence is an intrinsic feature of the design rather than a tunable choice.
From a design-based perspective, subsampling without replacement reveals that observation selection is not merely an implementation detail but a central mechanism through which random forests trade resolution against structural dependence.
6.5 Split-candidate restriction size $q$
Restricting the number of candidate variables considered at each split alters the dependence structure of the forest by changing how frequently independently generated trees induce the same local prediction rule at a prediction point $x$. This effect operates through alignment of the induced local averaging rules at $x$, as discussed in Section 4.3. In particular, restricting candidate variables reduces the opportunity for identical early split decisions along the path to $x$, thereby weakening alignment-induced dependence across trees.
Let $q\in\{1,\dots,p\}$ denote the number of candidate variables sampled at each internal node. Decreasing $q$ reduces the amount of shared candidate information available across trees at each split, which in turn reduces the probability that two trees make identical split decisions along the path to $x$. Here, two prediction rules are said to be identical at $x$ if they induce the same resampling-conditional weight vector $\bar{W}(x;\mathcal{I})$, equivalently if they define the same local averaging rule at $x$ after averaging over split-level randomization.
This mechanism can be made explicit through a simple combinatorial property of random subspace selection.

### v3

@@ 4.4 The covariance floor and its strict positivity
Figure 3: Empirical variance decomposition of $\hat{f}_{B}(x)$ as a function of the number of trees $B$ ($n=400$, $p=10$, $q=4$, bootstrap; DGM in Appendix B). A single covariate matrix $X$ is fixed and $R=300$ outcome vectors are drawn from the DGM. The total prediction variance (solid dark line) is the sample variance across draws, averaged over 20 test points; faint lines show individual test points before averaging. As $B$ increases, the Monte Carlo component (dashed line) goes to zero while the total variance converges to the mean covariance floor $\overline{C_{T}}=n_{\mathrm{pts}}^{-1}\sum_{k}C_{T}(x_{k})>0$ (horizontal line, shaded region), estimated independently from paired forests at B=8, [^0] B=8{,}000 trees. Agreement is within $1.3\%$, supporting the decomposition of Theorem 1.
@@
Sections 4.2 and 4.3 identified two distinct mechanisms contributing to $C_{T}(x)$: reuse of training observations in local averages at $x$, and alignment of the induced conditional prediction rules across trees. We now show that observation reuse alone is sufficient to guarantee strict positivity of the covariance floor, without invoking any alignment assumptions.
@@
Theorem 2 shows that repeated reuse of the same training outcomes in local averages at $x$ is sufficient to induce a strictly positive covariance floor. This effect is governed entirely by the tree-generating design acting on the fixed covariate configuration and persists regardless of the number of trees aggregated. The condition $\mathbb{P}\!\bigl(W_{i^{\star}}(x;\theta)>0\mid X\bigr)>0$ is satisfied by all standard implementations whenever $x$ lies in the support of the design and at least one nearby observation has positive conditional variance; moreover, the alignment component provides a second, independent nonnegative contribution to the covariance floor that is not required for strict positivity but reinforces it.
@@ 5.1 The infinite-aggregation target and resolution
Designs that assign expected weight to many observations are said to have lower resolution at $x$, while designs that concentrate expected weight on fewer observations near $x$ are said to have higher resolution at $x$.
@@
We use the term resolution rather than bias because it more precisely characterizes the fineness of the forest’s conditional prediction rule without invoking a parametric reference model. Coarser weight distributions average responses over larger sets of observations, yielding increased stability but reduced sensitivity to local variation in the outcomes. Finer weight distributions concentrate mass on fewer observations, producing predictors that adapt more sharply to local structure but that are more susceptible to dependence across trees through alignment of the induced conditional predictors. All forest design parameters affect resolution through their influence on the distribution of $W(x;\theta)$.
@@ 5.2 Training fraction and resolution scaling
A defining feature of random forest constructions is that each tree is trained on a strict subset of the available observations. Under bootstrap resampling, the expected number of distinct observations used to construct a tree is approximately $0.632n$ (Efron, 1983); under subsampling without replacement, the per-tree training size is fixed at $m=np_{\mathrm{obs}}$ for some training fraction $p_{\mathrm{obs}}\leq 1$. In both cases, the distribution of the weight vector $W(x;\theta)$ depends on $m$, and therefore the resolution of $f_{\infty}(x)$ is determined by the per-tree training size rather than by the full sample size.
@@
Whenever the distribution of $W(x;\theta)$ varies with the per-tree training size, $f_{\infty,m}(x)$ differs from $f_{\infty,n}(x)$ for $m<n$, and this discrepancy persists as the number of trees tends to infinity. Aggregation combines subsample-based conditional predictors but does not increase the resolution beyond that attainable by the underlying tree-level fits.
@@ 5.3 Structural error pathways
Resolution and dependence arise from different aspects of the same design-induced randomness under fixed $X$. The infinite-aggregation predictor $f_{\infty}(x)=\mathbb{E}[T_{\theta}(x)\mid X]$ is determined by the average weight vector $\mathbb{E}[W(x;\theta)\mid X]$, which governs the resolution of the forest at $x$. The covariance floor $C_{T}(x)$ depends on the variability and co-movement of the random weights across design realizations. Any design choice that reduces dependence necessarily alters the distribution of these weight vectors, and therefore also affects the resolution of the infinite-aggregation predictor. This relationship is analogous to the classical bias–variance tradeoff: the forest design redistributes error between resolution and dependence, and no choice of design parameters can optimize both simultaneously.
@@ 6.1 Parameterizing the tree-generating mechanism
We parameterize the tree-generating mechanism $\theta$ through a collection of design components (i.e. hyperparameters) defined relative to the observed sample size $n$. Although $n$ is not a tuning parameter in the same sense as subsampling or split restriction, it plays a structural role by determining the support over which the induced averaging sets $A_{\theta}(x)$ and weight vectors $W(x;\theta)$ are defined, and therefore influences both resolution and dependence.
@@
    Observation selection. A random index set $\mathcal{S}_{\theta}\subset\{1,\dots,n\}$ is drawn with fixed cardinality $m=|\mathcal{S}_{\theta}|$, defining the per-tree observation-selection fraction $p_{\mathrm{obs}}=m/n$. The corresponding indicator vector $\mathcal{I}_{\theta}=(I_{\theta,1},\dots,I_{\theta,n})$, where $I_{\theta,i}:=\mathbf{1}\{i\in\mathcal{S}_{\theta}\}$, encodes which observations are exposed to the tree construction.
@@
    Split-candidate restriction. At each internal node, a random candidate set of variables $\mathcal{J}_{\theta,v}\subset\{1,\dots,p\}$ is drawn with fixed size $q$. The split at that node is chosen by optimizing the splitting criterion over $\mathcal{J}_{\theta,v}$.
@@
    Terminal occupancy constraint. A stopping rule enforces a lower bound $s\geq 1$ on the size of the induced averaging set, $|A_{\theta}(x)|\geq s$ almost surely, ensuring that each tree-level prediction averages at least $s$ observed outcomes. A maximum tree depth $d_{\max}$ operates as an alternative constraint on partition fineness by limiting the number of recursive splits along any root-to-leaf path.
@@
    Aggregation level. The forest predictor averages $B$ independent realizations $\theta_{1},\dots,\theta_{B}$. This parameter governs only Monte Carlo approximation error: it does not affect the distribution of the underlying tree-level conditional predictor $T_{\theta}(x)$, and therefore has no effect on resolution, dependence, or the covariance floor.
@@
The forest design is therefore characterized by the parameter vector $(n,p_{\mathrm{obs}},q,s,B)$. For fixed $n$, the parameters $(p_{\mathrm{obs}},q,s)$ jointly determine the distribution of the induced averaging sets $A_{\theta}(x)$ and weight vectors $W(x;\theta)$, and thereby control both the resolution of the infinite-aggregation target and the dependence structure across trees. The role of $B$ is purely to approximate this design-based target through Monte Carlo averaging. The following subsections analyze how each remaining structural design parameter acts on resolution and dependence.
6.2 Observation-selection fraction $\boldsymbol{p}_{\mathrm{obs}}$
@@
The observation-selection fraction $p_{\mathrm{obs}}$ directly alters both the dependence structure across trees and the resolution of the induced predictor. From the perspective of dependence, $p_{\mathrm{obs}}$ acts through the observation-overlap component of the covariance decomposition via joint inclusion probabilities
@@
which determine how frequently the same observation can receive positive weight at $x$ in multiple trees. Increasing $p_{\mathrm{obs}}$ increases these probabilities and, holding the within-tree partitioning mechanism fixed, strengthens the observation-overlap contribution, thus increasing $C_{T}(x)$.
Beyond its effect on dependence, $p_{\mathrm{obs}}$ also governs resolution by restricting which observations are available to define candidate splits and terminal regions. Smaller values of $p_{\mathrm{obs}}$ limit the data available to each tree, inducing coarser averaging and lower resolution. Larger values increase resolution at the cost of stronger reuse-driven dependence between trees. Subsampling without replacement makes this trade-off explicit: $p_{\mathrm{obs}}$ becomes a directly tunable parameter controlling both resolution and dependence.
The bootstrap corresponds to a particular, non-tunable choice of $p_{\mathrm{obs}}$. Although each tree is grown from a resample of size $n$, the expected fraction of distinct observations satisfies
$\mathbb{P}(i\in\mathcal{I}_{\theta})=1-(1-1/n)^{n}\approx 1-e^{-1},\qquad\mathbb{P}(i\in\mathcal{I}_{\theta}\cap\mathcal{I}_{\theta^{\prime}})\approx(1-e^{-1})^{2},$
so the observation-selection fraction and the resulting magnitude of observation-induced dependence are intrinsic features of the bootstrap design rather than tunable choices.
6.3 Split-candidate restriction size $\boldsymbol{\mathrm{(}q)}$
Restricting the number of candidate variables at each split influences the dependence between trees by changing how frequently independently generated trees make similar splitting decisions along the path to $x$. This effect operates through the alignment component of the covariance decomposition.
The mechanism can be made explicit through a combinatorial property of how candidate variables are sampled at each split.